OpenStudy (anonymous):

if f(x)=1/x, find f(x+h)

OpenStudy (anonymous):

The meaning of f(x) = 1/x is that for everything inside the parentheses, you plop that expression right into everywhere you see x. The notation is confusing but just redefine it as f(z) = 1/z. There's nothing wrong with that. so now evaluate f(z) when z = x+h. So, f(x+h) = 1/(x+h), simple as that.

OpenStudy (anonymous):

Thank you.